The owners' equity accounts for Vidi International are shown here: Common stock ($.50 par value) $ 42,500 Capital surplus 355,000 Retained earnings 778,120 Total owners’ equity $ 1,175,620 a-1. If the company declares a five-for-one stock split, how many shares are outstanding now? (Do not round intermediate calculations.) a-2. What is the new par value per share? (Do not round intermediate calculations and round your answer to 3 decimal places, e.g., 32.161.) b-1. If the company declares a one-for-four reverse stock split, how many shares are outstanding now? (Do not round intermediate calculations.) b-2. What is the new par value per share?
Answer #1

Number of shares outstanding currently = 42,500/0.5 = 85,000 shares

A five for one stock split means that 5 shares will be issued in place of every one share

a-1 Hence, shares outstanding now = 85,000*5 = 425,000

a-2 New Par value per share = 0.50/5 = $0.10

One for Four Reverse Stock Split means that one share will be issued in place of every 4 shares

b-1 Number of shares outstanding = 425,000/4 = 106,250 shares

b-2 Par Value per share = 0.10*4 = $0.40
Answer #2

a-1. If the company declares a five-for-one stock split, the number of shares outstanding will increase fivefold. To find the new number of shares, we divide the current total owners' equity by the new par value per share.

Current number of shares = Total owners’ equity / Current par value per share

Total owners’ equity after the split = $1,175,620 (Given) New par value per share = Current par value per share / 5 = $0.50 / 5 = $0.10

New number of shares = Total owners’ equity after the split / New par value per share

New number of shares = $1,175,620 / $0.10 New number of shares = 11,756,200 shares

a-2. The new par value per share after the five-for-one stock split is $0.10.

b-1. If the company declares a one-for-four reverse stock split, the number of shares outstanding will decrease. To find the new number of shares, we divide the current total owners' equity by the new par value per share.

Current number of shares = Total owners’ equity / Current par value per share

Total owners’ equity after the reverse split = $1,175,620 (Given) New par value per share = Current par value per share * 4 = $0.50 * 4 = $2.00

New number of shares = Total owners’ equity after the reverse split / New par value per share

New number of shares = $1,175,620 / $2.00 New number of shares = 587,810 shares

b-2. The new par value per share after the one-for-four reverse stock split is $2.000 (rounded to 3 decimal places).

answered by: Hydra Master